### v1

# Effective and Transparent RAG: Adaptive-Reward Reinforcement Learning for Decision Traceability

Abstract
Retrieval-Augmented Generation (RAG) has significantly improved the performance of large language models (LLMs) on knowledge-intensive domains. However, although RAG achieved successes across distinct domains, there are still some unsolved challenges: 1) Effectiveness. Existing research mainly focuses on developing more powerful RAG retrievers, but how to enhance the generator’s (LLM’s) ability to utilize the retrieved information for reasoning and generation? 2) Transparency. Most RAG methods ignore which retrieved content actually contributes to the reasoning process, resulting in a lack of interpretability and visibility. To address this, we propose ARENA (Adaptive-Rewarded Evidence Navigation Agent), a transparent RAG generator framework trained via reinforcement learning (RL) with our proposed rewards. Based on the structured generation and adaptive reward calculation, our RL-based training enables the model to identify key evidence, perform structured reasoning, and generate answers with interpretable decision traces. Applied to Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct, abundant experiments with various RAG baselines demonstrate that our model achieves 10–30% improvements on all multi-hop QA datasets, which is comparable with the SOTA Commercially-developed LLMs (e.g., OpenAI-o1, DeepSeek-R1). Further analyses show that ARENA has strong flexibility to be adopted on new datasets without extra training. Our models and codes are publicly released [^1] https://github.com/ren258/ARENA.
1 Introduction
Retrieval-Augmented Generation (RAG) has become a powerful paradigm for enhancing large language models (LLMs) with non-parametric knowledge [8, 12, 25, 55], particularly in knowledge-intensive domains such as medicine [23], law [49], and finance [37]. While retrieved information can significantly improve generation performance [22, 52], existing research has mainly concentrated on improving retrievers, with relatively less attention given to the generator component. Our empirical analysis reveals a notable performance gap: on widely-used and challenging multi-hop QA benchmarks, 7B-scale models using identical retrieved contexts underperform state-of-the-art reasoning models like OpenAI-o1 [30] and DeepSeek-R1 [42] by 15–35% in answer accuracy (see Appendix A). This indicates that the generator’s reasoning capability, not just the quality of retrieval, remains a key bottleneck in current RAG pipelines.
Beyond boosting effectiveness, advancing the generator opens up valuable opportunities for improving transparency — critical qualities for deploying LLMs in real-world, high-stakes applications to understand which retrieved document is involved in reasoning and generation [41, 48, 24]. Many generators produce unstructured answers without exposing the underlying decision process, limiting their interpretability and trustworthiness. The central challenge lies in enabling the generator to identify, integrate, and reason over relevant evidence from noisy or redundant contexts [13]. Smaller open-source models face greater difficulty in such scenarios due to limited reasoning capacity [9].
Recently, reinforcement learning (RL) has shown promise in improving the reasoning ability of LLMs, especially in math [53] and code [26] generation, which often rely on reward signals derived from structured task objectives, enabling models to learn reasoning behavior beyond what supervised fine-tuning can offer. This inspires us to explore RL as a path for improving RAG generation. Existing RL-based approaches—such as GRPO [11]—typically use generic reward designs and output formats, which fail to capture the structure of multi-hop QA [34]. Moreover, unstable KL regularization often leads to divergence during training, limiting their applicability in retrieval-grounded tasks.
To address these challenges, we propose ARENA (Adaptive-Rewarded Evidence Navigation Agent), a transparent and effective RAG generator framework trained via reinforcement learning. ARENA introduces a structured output format that includes selected references, explicit reasoning traces, and final answers, enabling end-to-end interpretability. We further design a suite of adaptive, task-specific rewards and a stabilized optimization strategy tailored for multi-hop QA. Applied to open-source models like Qwen2.5-7B-Instruct [50] and Llama3.1-8B-Instruct [10], ARENA achieves 10–30% accuracy improvements across three datasets, rivaling the best closed-source systems while offering better transparency and generalizability. Our main contributions are listed as follows:
  • We identify two key challenges in current RAG systems, namely limited reasoning ability and lack of interpretability, and call for greater attention to the generator optimization.
  • We propose ARENA, a reinforcement learning-based framework for RAG that integrates structured generation, adaptive reward design, and stabilized training for transparent and effective RAG reasoning.
  • Experiments demonstrate that ARENA outperforms all open-source baselines of comparable scale, achieving results on par with leading closed-source models across three challenging multi-hop QA datasets. ARENA also generalizes well across datasets and backbones.
  • All the codes and models are released to support future study on optimizing RAG generator.
2 Related Work
2.1 Retrieval-Augmented Generation
RAG has emerged as an effective approach to enhancing model performance across a wide range of NLP tasks by incorporating externally retrieved documents into the input [14, 8], particularly for knowledge-intensive applications such as open-domain question answering (ODQA) [22]. By grounding the model’s generation in retrieved evidence, RAG can substantially mitigate hallucinations and factual inaccuracies [40, 2]. Moreover, recent studies have demonstrated that the benefits of retrieval become increasingly pronounced as model scale grows [28, 39], further underscoring the practical value of retrieval-based methods. In the ODQA task, the prevailing paradigm follows a retriever-and-reader framework [1], where significant effort has been devoted to enhancing the retriever. Approaches include decomposing complex queries into sub-questions using prompts [33], rewriting queries to enhance retrieval quality [27], guiding retrieval through CoT reasoning [46], constructing summaries to select the most plausible answer among candidates [20], and compressing prompts while integrating retrieval information [18]. Despite these substantial advances, how to optimize the RAG generator was ignored by most studies.
The multi-hop task [51, 15, 45, 33] poses significant challenges in the field of RAG, as they require LLMs to integrate information across multiple retrieved documents in order to generate a correct answer, which requires more powerful reasoning and generation ability. However, the presence of irrelevant or noisy contexts can adversely affect answer generation, resulting in factual errors [32, 6]. Thus, strengthening the reader’s capability to reason over and selectively utilize retrieved content is critical for achieving robust performance.
Beside, some studies have demonstrated that verifying the reliability of LLM-generated content based on citations is an effective approach to mitigating hallucinations and provides transparent RAG for users [17]. But most of these studies are post-hoc based, which means the citations are not generated during LLM reasoning. We propose that, this step can also be finished by the RAG generator.
2.2 LLMs with RL
RL serves as a powerful tool for LLMs to improve their alignment to human values and reasoning abilities. In particular, reinforcement learning with human feedback (RLHF) has demonstrated the effectiveness of aligning model outputs with human preferences [31]. Beyond alignment, models such as OpenAI-o1 [30], DeepSeek-R1 [11] and Kimi k1.5 [44] showcase significant gains in mathematical reasoning and code generation following RL-based training. More recently, R1-Searcher [42] and ReSearch [4] have leveraged Reinforce++ [16] and GRPO [38], respectively, to enhance the abilities of the models to interact with retrieval systems and produce accurate responses. Collectively, these methods highlight how reinforcement learning reward signals can systematically improve the performance of LLM in unsupervised settings. However, prior methods have primarily focused on strengthening general reasoning ability, without explicitly designing reward functions tailored to specific tasks such as RAG. In this work, we introduce specialized incentives to address this gap.
Figure 1: Overview of the ARENA (Adaptive-Rewarded Evidence Navigation Agent) framework. The system is composed of three key components: ① Structured Generation, where the policy model generates a multi-part response including selected evidence, reasoning analysis, and a final answer; ② KL Stabilization, we modify the KL formulation to improve stability; ③ Adaptive Reward Calculation, where the model outputs are evaluated on four axes—format, accuracy, relevance, and bonus—to provide interpretable and fine-grained training signals.
3 ARENA: Adaptive-Rewarded Evidence Navigation Agent
We introduce ARENA (Adaptive-Rewarded Evidence Navigation Agent), a novel reinforcement learning framework for the training of RAG generators. ARENA enhances the reasoning capabilities of language models and improves decision transparency, enabling more interpretable and reliable deployment in knowledge-intensive domains. As shown in Figure 1, ARENA consists of three key components: (1) Structured generation, where the model produces evidence-grounded multi-part responses; (2) KL stabilization, we stabilized KL regularization for policy refinement; and (3) Adaptive reward calculation, which provides interpretable feedback across multiple dimensions. We will introduce these modules in detail in the following subsections.
3.1 Framework formalization
Given a natural language question $q$ and a set of $k$ retrieved reference passages $\mathcal{C}=\{c_{1},c_{2},...,c_{k}\}$, the ARENA generator produces a structured response with three clearly defined components:
  • A set of selected reference indices $\mathcal{I}\subseteq\{1,2,...,k\}$ indicating which passages are used.
  • A reasoning trace $\mathcal{Z}$ that synthesizes the selected passages to generate the final answer.
  • A concise final answer $\mathcal{O}$ derived from the reasoning trace.
This defines the generator as a mapping:
$(q,\mathcal{C})\mapsto(\mathcal{I},\mathcal{Z},\mathcal{O})$
To support training and evaluation, we construct a structured prompt format as shown in Table 1 (Figure 1 Module 1). Notably, we introduce an explicit <relevance> part to enforce interpretable evidence selection, which is critical for transparent multi-hop reasoning. This part requires the model to generate only the reference indices it relies on, making the reasoning process externally verifiable.
Table 1: ARENA prompt template format. The <relevance> part is the key to encourage explicit and auditable evidence selection. The placeholders {question} and {references} are filled in at runtime.
A conversation between User and Assistant. The user asks a question and gives some references. The assistant should answer the question based on the references.
User’s input will always contain:
<question>[ the question to answer ]</question>
<references>[ references starting with numbers ]</references>
Assistant’s response must contain EXACTLY three sections:
<relevance> [list ONLY reference numbers that provide useful information in square brackets, e.g. [1,5]]</relevance>
<analysis>[ combine information from relevant references to build the answer. Explicitly mention which references support each claim ]</analysis>
<answer>[ answer with ONLY a short phrase or single word. no explanations ]</answer>
User:
<question>{question}</question>
<references>{references}</references>
Each generation component serves a specific role/function: the <relevance> part ensures explicit evidence grounding, the <analysis> part compels the LLM to articulate its reasoning based on the selected references, and the <answer> part provides a concise and definitive result. Together, this format enforces traceable generation and supports external auditing, aligning with ARENA’s goals of enhancing reasoning quality and interpretability.
3.2 Reinforcement learning with stability enhancements
In complex reasoning tasks, supervised fine-tuning (SFT) often falls short due to the lack of high-quality, annotated reasoning traces [5]. Recently, reinforcement learning (RL) has emerged as a powerful alternative to enhanceperformance through self-improvement. Motivated by this, ARENA employs reinforcement learning to activate and refine the reasoning capabilities of instruction-tuned models to enhance the RAG generation ablities of them.
Brief review of Group Relative Policy Optimization (GRPO)
GRPO [38] is an advanced policy optimization algorithm that improves over Proximal Policy Optimization (PPO) [36] by eliminating the need for a critic model. For each training instance $q$, the policy $\pi_{\theta}$ generates $G$ outputs $\{o_{i}\}_{i=1}^{G}$. Each output is scored using the task-specific reward function (detailed in Section 3.3), and a group-wise normalized advantage is computed as: $\mathcal{A}_{i}=\min\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},\text{clip}\left(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},1-\epsilon,1+\epsilon\right)\right)\frac{r_{i}-{\mathrm{m}ean(\{r_{1},r_{2},\cdots,r_{G}\})}}{{\mathrm{s}td(\{r_{1},r_{2},\cdots,r_{G}\})}}$, where $r_{i}=\sum_{j=1}^{N}w_{j}\cdot R_{j}(o_{i}\mid q)$ is the weighted reward for output $o_{i}$.
The KL divergence between the current policy and a fixed reference model is incorporated as a regularization term to prevent policy collapse. The final GRPO objective is:
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{q,\{o_{i}\}\sim\pi_{\theta_{old}}}\left[\frac{1}{G}\sum_{i=1}^{G}\left(\mathcal{A}_{i}-\beta\,\mathbb{D}_{KL}\left(\pi_{\theta}\|\pi_{ref}\right)\right)\right].$
KL divergence stabilization
The KL divergence term plays a key role in constraining the updated policy from diverging excessively from the reference model. In standard GRPO, it is estimated using the following unbiased formulation:
$r=\frac{\pi_{\mathrm{ref}}(o\mid q)}{\pi_{\theta}(o\mid q)},\quad\mathbb{D}_{KL}(\pi_{\theta}\|\pi_{\mathrm{ref}})=r-\log r-1,$
which is theoretically appealing: it is non-negative, unbiased, and has low variance under common assumptions. However, this estimator becomes numerically unstable when the policy assigns near-zero probability to sampled outputs ($\pi_{\theta}(o\mid q)\to 0$), causing the importance ratio $r\to\infty$. This results in gradient explosions and erratic updates that severely degrade training stability.
To address this, we aim to retain the desirable properties of the original KL estimator—namely, non-negativity, low variance, and smooth penalization—while ensuring numerical stability in rare-event sampling. Motivated by this, we use a simple but effective replacement [35]:
$\text{KL}_{\mathrm{stable}}(o)=\frac{1}{2}(\log r)^{2},$
which maintains non-negativity and provides a symmetric, smooth penalty for large divergences. It avoids singularities even when $\pi_{\theta}(o\mid q)$ is small, thus offering more stable gradients in training.
3.3 Adaptive reward design
While recent math-oriented studies demonstrated that simple format and accuracy rewards suffice to activate reasoning capabilities in GRPO, we find that such task-agnostic metrics fall short in capturing the nuanced objectives of RAG generation. To this end, we propose a set of task-specific, interpretable reward functions that provide fine-grained supervision aligned with reasoning quality and evidence traceability, namely:
Format Reward $R_{\mathrm{format}}(o_{i}\mid q)$.
Outputs that match the expected structural format—consisting of <relevance>, <analysis>, and <answer> parts in order—receive a reward of 1; others receive 0.
Accuracy Reward $R_{\mathrm{accuracy}}(o_{i}\mid q)$.
To evaluate the correctness of the final answer, we apply normalized Exact Match. The output from the <answer> field is lowercased, stripped of punctuation and articles, and compared to the gold answer. If the normalized strings match exactly, the model receives a reward of 1; otherwise 0. Exact Match is more effective than F1/LLM judgment here.
Relevance Reward $R_{\mathrm{relevance}}(o_{i}\mid q)$.
This reward measures whether the model correctly identifies the supporting evidence. The predicted reference indices from the <relevance> section are compared with ground truth. A full match yields 1 point, partial overlap yields 0.5, and no overlap yields 0. This encourages models to explicitly ground their reasoning in relevant sources.
Bonus Reward $R_{\mathrm{bonus}}(o_{i}\mid q)$.
To promote holistic reasoning behavior, we add a high-value bonus reward when the model simultaneously satisfies all three criteria above. If the format, accuracy, and relevance rewards are all 1, an additional bonus of 10 is assigned; otherwise, the bonus is 0. This reinforces complete, well-aligned outputs.
The final reward used to train the policy model is computed as the sum of the four components and no reweighting strategy/training is adopted:
$r_{i}=R_{\mathrm{format}}(o_{i}\mid q)+R_{\mathrm{accuracy}}(o_{i}\mid q)+R_{\mathrm{relevance}}(o_{i}\mid q)+R_{\mathrm{bonus}}(o_{i}\mid q),$
where all weights are set to 1 by default. This combined signal balances structural, semantic, and evidential quality in a unified reward framework.
4 Experimental Setups
4.1 Datasets
We use three widely-used and challenging multi-hop QA datasets for evaluation: HotpotQA [51], 2WikiMultiHopQA [15], and MuSiQue [45]. Each QA consists of a natural language question, a final answer, and 10–20 retrieved paragraphs from Wikipedia (predefined in the public datasets). We randomly selected a portion from the whole dataset for training and testing as previous studies [42, 47, 19], statistics are shown in Table 2. Detailed dataset preprocessing steps are described in Appendix B.
Table 2: Detailed dataset statistics. Hop counts are derived from supporting_facts fields.
Dataset Split Data Size # Paragraphs 1-hop 2-hop 3-hop 4-hop Avg. Hops
HotpotQA Train 10,000 10 3 9977 20 0 2.00
HotpotQA Test 500 10 0 500 0 0 2.00
2WikiMultiHopQA Train 10,000 10 7 7839 3 2151 2.39
2WikiMultiHopQA Test 500 10 0 406 0 94 2.19
MuSiQue Train 5,000 20 0 3595 1107 298 2.85
MuSiQue Test 500 20 0 251 153 96 2.73
All Train 25,000 10 21,411 1130 2449 2.30
4.2 Evaluation metrics
Following prior work [51, 8], we evaluate model performance using three metrics: normalized Exact Match (EM), F1 score (F1), and LLM-as-a-Judge (LJ) [54]. EM and F1 capture surface-level correctness, while LJ uses GPT-4o [29] (version 2024-11-20) to assess semantic correctness. The implementation details of all three metrics are provided in Appendix C.
4.3 Baselines
Prompt-based RAG models. We select open-source reasoning models of similar scale, including DeepSeek-R1-Distill-Qwen-7B and Qwen3-8B, as well as our backbone models Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct. Those models answer the question with references directly [^2] For reasoning-capable models, the internal thinking process is implicitly conducted during generation in the direct setting.. We also include two prompting-based methods: SuRe [20], which enhances LLM performance via summarized retrieval, and self-ask [33], which decomposes complex questions into sub-questions and answers them iteratively. These serve as strong zero-shot and few-shot baselines.
SFT-based RAG models. We include two supervised fine-tuning baselines: SFT-direct, which directly generates the final answer, and SFT-reasoning, a reasoning-capable model that outputs structured responses. Training details for both models are provided in Appendix D.
RL-based RAG models. This category includes models trained with reinforcement learning. We first implement a Naive GRPO baseline using the original GRPO algorithm [38], where models are trained to produce <think></think> and <answer></answer> format responses with format and accuracy reward. Three SOTA RL-based reasoning frameworks are also involved: R1-Searcher [42], SimpleDeepSearcher [43], and ReSearch [3]. Each of these methods introduces a unique approach to reasoning—such as external search invocation, distillation-driven retrieval, or search-integrated generation. For consistency, we use their released checkpoints and replace their retrieved contexts with those provided by our datasets.
Table 3: Main experimental results across three datasets and three evaluation metrics. Models are grouped by training methodology. Our ARENA models are placed alongside their corresponding base models at the bottom for direct comparison. Bold indicates the best score; underline indicates the second-best. Metrics: EM = Exact Match (%), F1 = F1 score (%), LJ = LLM-as-a-Judge (%).
Model HotpotQA 2WikiMultiHopQA MuSiQue
EM F1 LJ EM F1 LJ EM F1 LJ
Prompt-based models
DeepSeek-R1-Distill-Qwen-7B 0.332 0.487 0.712 0.290 0.407 0.658 0.116 0.184 0.278
Qwen3-8B 0.582 0.719 0.768 0.652 0.727 0.784 0.336 0.397 0.394
SuRe-GPT-4o 0.490 0.695 0.742 0.490 0.607 0.632 0.194 0.307 0.324
SuRe-Qwen-7B 0.504 0.648 0.696 0.364 0.448 0.440 0.146 0.233 0.226
self-ask-GPT-4o 0.408 0.577 0.662 0.430 0.565 0.614 0.162 0.305 0.346
self-ask-Qwen-7B 0.016 0.112 0.246 0.004 0.128 0.086 0.006 0.052 0.087
SFT-based models
Qwen-7B-SFT-direct 0.498 0.628 0.662 0.540 0.609 0.620 0.168 0.254 0.214
Qwen-7B-SFT-reasoning 0.400 0.532 0.599 0.536 0.615 0.622 0.112 0.172 0.174
RL-based models
Naive-GRPO 0.534 0.662 0.732 0.620 0.684 0.720 0.332 0.435 0.420
R1-Searcher-7B 0.590 0.731 0.791 0.632 0.704 0.730 0.250 0.370 0.327
SimpleDeepSearcher-7B 0.500 0.633 0.686 0.632 0.707 0.748 0.258 0.347 0.344
ReSearch-7B 0.578 0.729 0.778 0.586 0.658 0.684 0.300 0.398 0.407
Ours (ARENA)
Llama3.1-8B-Instruct (Original) 0.528 0.676 0.738 0.398 0.478 0.484 0.248 0.370 0.324
ARENA-Llama-8B 0.552 0.703 0.780 0.622 0.710 0.735 0.358 0.445 0.466
Qwen2.5-7B-Instruct (Original) 0.484 0.628 0.660 0.334 0.424 0.412 0.252 0.354 0.306
ARENA-Qwen-7B 0.628 0.762 0.812 0.660 0.752 0.786 0.400 0.520 0.508
4.4 Implementation details
We conduct ARENA on Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct. Our framework is implemented based on the Open-R1 [7]. During inference, we concatenate all retrieved paragraphs provided by the dataset in their original order. All models are trained using 8 NVIDIA A100-80G GPUs, with 7 GPUs allocated for policy optimization and 1 GPU dedicated to rollout inference via a vLLM [21] engine. More training details are introduced in Appendix E.
5 Experimental Results
5.1 Main results
As shown in Table 3, ARENA-Qwen-7B achieves the best performance across all datasets and metrics, consistently outperforming all baselines of similar scale. Compared to its backbone models, ARENA brings 10–30% improvements, with particularly strong gains on MuSiQue, highlighting its ability to handle longer and more complex reasoning chains. ARENA-Llama-8B also outperforms the original Llama3.1-8B-Instruct significantly, showing that our method is capable of distinct backbones. We also observe that RL-based models consistently outperform SFT-based models, confirming the advantage of learning through reward-guided exploration. In contrast, supervised fine-tuning fails to deliver meaningful gains, suggesting that it struggles to equip models with structured reasoning.
Table 4: Evaluation of ARENA-Qwen-7B under different training dataset combinations. “+WM” denotes training on 2WikiMultiHopQA + MuSiQue, “+HM” on HotpotQA + MuSiQue, and “+HWM” on all three datasets. Results are reported on all three test sets to assess cross-domain generalization.
Model HotpotQA 2WikiMultiHopQA MuSiQue
EM F1 LJ EM F1 LJ EM F1 LJ
ARENA-Qwen-7B + WM 0.568 0.711 0.764 0.642 0.732 0.734 0.394 0.504 0.486
ARENA-Qwen-7B + HM 0.598 0.741 0.796 0.578 0.671 0.680 0.404 0.519 0.478
ARENA-Qwen-7B + HWM 0.628 0.762 0.812 0.660 0.752 0.774 0.400 0.520 0.508
5.2 Generalization ability
The generalization ability of ARENA-Qwen-7B when trained on different subsets of the available datasets is shown in Table 4. Specifically, “+WM” and “+HM” denote training on two datasets while holding out one, allowing us to assess zero-shot transfer to an unseen domain. In both cases, the model performs strongly on the held-out dataset, indicating that the reasoning capabilities learned by ARENA generalize well across different data distributions. This supports our claim that the improvements from our method are not tied to a specific dataset, but reflect transferable multi-hop reasoning skills.
Figure 2: Ablation results comparing Naive GRPO, ARENA without KL stabilization, and full ARENA on three evaluation metrics across datasets. ARENA shows consistent improvements in output quality and interpretability.
Figure 3: A Case Study of Naive GRPO and ARENA on the HotpotQA. ARENA correctly extracts relevant paragraphs, provides a structured explanation, and delivers the right answer.
5.3 Ablation study
We evaluate the effect of ARENA’s key components by comparing three variants: Naive GRPO (basic reward and flat output format), w/o KL stabilization (adaptive reward without stable KL estimation), and the full ARENA-Qwen-7B. As shown in Figure 2, ARENA achieves the best results across all datasets in LLM-as-a-Judge score, format score, and relevance score, demonstrating its more powerful ability in generate accurate and interpretable outputs.
Table 5: Comparison between ARENA-Qwen-7B and three strong closed-source models on HotpotQA. ARENA shows strong competitiveness despite smaller model size.
Model EM F1 LJ
GPT-4o 0.628 0.788 0.820
OpenAI-o1 0.656 0.816 0.890
DeepSeek-R1 0.618 0.781 0.836
ARENA-Qwen-7B 0.628 0.762 0.812
Table 6: Comparison of different accuracy reward formulations used during RL training. EM-based reward leads to the best performance on HotpotQA across multiple evaluation dimensions.
Judge Method EM F1 LJ
GPT-4o 0.612 0.752 0.804
F1 0.600 0.738 0.794
EM (ours) 0.628 0.762 0.812
5.4 Further analyses
ARENA Vs. SOTA Commercially-developed Models. We compare ARENA-Qwen-7B with three strong commercially-developed LLMs on HotpotQA in subsection 5.3, including GPT-4o &\&& OpenAI-o1 from OpenAI and DeepSeek-R1. Despite its smaller size, our ARENA shows competitive results across all metrics, with performance close to GPT-4o in particular. This highlights ARENA’s potential as a lightweight yet effective alternative for interpretable multi-hop reasoning.
Reward Analysis. We also test different reward formulations for the accuracy component during RL training, including EM, F1, and LLM-based judgment. As shown in subsection 5.3, using Exact Match as the reward provides the best overall results across EM, F1, and LJ evaluation. This confirms that a simple, binary EM-based reward provides a more stable and effective training signal for multi-hop QA tasks. More experimental results and analyses on 2WikiMultiHopQA and MuSiQue datasets are provided in Appendix G.
Case Study. Figure 3 illustrates a representative case. While Naive GRPO produces a valid-looking but unsupported answer, ARENA identifies the correct references and generates a transparent reasoning path. The benefit of KL stabilization is further supported by training curves (see Appendix F), where the stabilized version maintains smooth KL values and avoids extreme divergence spikes.
6 Conclusion
We propose ARENA, an Adaptive-Rewarded Evidence Navigation Agent that enhances the generator component of RAG systems via reinforcement learning with task-specific rewards. ARENA enables models to identify relevant evidence, reason in a structured and interpretable way, and produce accurate, verifiable answers. Extensive experiments across three multi-hop QA benchmarks demonstrate that ARENA significantly outperforms all open-source models of similar scale, with 10–30% gains in accuracy. It also generalizes well across datasets and backbone models, highlighting its robustness and transferability. By generating explicitly structured outputs, ARENA facilitates transparent reasoning, addressing a critical need for interpretability in real-world applications.
Limitation
Despite its effectiveness and transparent, ARENA assumes access to high-quality retrieved content and may underperform when critical evidence is missing or noisy. Its reward and KL stabilization are designed for RAG, which may unable to be direct adopted to other domains. Future work may explore retrieval-aware reward shaping and extend the framework to other tasks.
Appendix A Settings of the Pilot Experiments
We present a comprehensive comparison of multiple vanilla models under two inference settings: direct, where the model generates the answer based only on the question, and direct RAG, where the model receives both the question and all retrieved contexts as input. As shown in Table 7, RAG inputs consistently improve the performance of all models, highlighting the effectiveness of retrieval-augmented generation.
Moreover, even under the same RAG context, different LLMs exhibit large performance discrepancies. Notably, models like OpenAI-o1 and DeepSeek-R1 outperform open-source 7B models by 15–35% across multiple datasets and metrics. These results underscore the variability of RAG generator capabilities and motivate our work to improve them through reinforcement learning.
Table 7: Performance of vanilla models under two inference strategies: direct and direct RAG. All models are evaluated using identical retrieved contexts. Bold indicates the highest value per column. Metrics: EM = Exact Match (%), F1 = F1 score (%), LJ = LLM-as-a-Judge (%).
Model HotpotQA 2WikiMultiHopQA MuSiQue
EM F1 LJ EM F1 LJ EM F1 LJ
direct
Qwen2.5-7B-Instruct 0.188 0.247 0.262 0.222 0.253 0.234 0.040 0.090 0.061
Llama3.1-8B-Instruct 0.194 0.283 0.284 0.210 0.246 0.232 0.038 0.080 0.054
DeepSeek-R1-Distill-Qwen-7B 0.086 0.150 0.162 0.144 0.206 0.234 0.012 0.043 0.026
Qwen3-8B 0.248 0.344 0.367 0.220 0.263 0.264 0.030 0.092 0.060
GPT-4o 0.378 0.500 0.500 0.342 0.386 0.366 0.100 0.192 0.124
OpenAI-o1 0.510 0.667 0.726 0.530 0.634 0.684 0.292 0.404 0.392
DeepSeek-R1 0.422 0.564 0.590 0.454 0.536 0.546 0.202 0.330 0.294
direct RAG
Qwen2.5-7B-Instruct 0.484 0.628 0.660 0.334 0.424 0.412 0.252 0.354 0.306
Llama3.1-8B-Instruct 0.528 0.676 0.738 0.398 0.478 0.484 0.248 0.370 0.324
DeepSeek-R1-Distill-Qwen-7B 0.332 0.487 0.712 0.290 0.407 0.658 0.116 0.184 0.278
Qwen3-8B 0.582 0.719 0.768 0.652 0.727 0.784 0.336 0.397 0.394
GPT-4o 0.628 0.788 0.820 0.606 0.687 0.717 0.505 0.617 0.593
OpenAI-o1 0.656 0.816 0.890 0.702 0.805 0.837 0.622 0.744 0.762
DeepSeek-R1 0.618 0.781 0.836 0.721 0.801 0.838 0.559 0.702 0.703
Appendix B Dataset Preprocessing
We follow the preprocessing pipeline from PIKE-RAG [47] to format the train and dev splits for all datasets. Each example includes:
  • A natural language question.
  • A list of retrieved paragraphs from Wikipedia.
  • A gold-standard answer (often as a list of acceptable synonyms).
  • A set of supporting_facts, which mark the sentences essential for answering the question.
To construct the <references> field, we concatenate all retrieved paragraphs in their original order with paragraph-level IDs (1 to $n$). To supervise relevance prediction and reasoning, we use the supporting_facts annotations to locate which paragraphs contain the necessary sentences. These are mapped to their paragraph indices to produce a ground-truth list of relevant reference IDs.
For training, we randomly sample 10,000 instances each from HotpotQA and 2WikiMultiHopQA, and 5,000 from MuSiQue. For testing, we sample 500 development examples from each dataset, following the setup in [42, 47, 19] to ensure comparability.
Answer Format.
The answer labels are often represented as a list of acceptable values (e.g., synonyms, numbers, yes/no), and models are considered correct if their normalized output matches any item in the list. This setting is used consistently across all automatic and LLM-based evaluations.
Appendix C Details of Evaluation Metrics
We use three evaluation metrics in our experiments: Exact Match (EM), F1 Score (F1), and LLM-as-a-Judge (LJ). Below we describe the implementation details of each.
C.1 Exact match and normalization
The Exact Match (EM) metric determines whether the model’s answer exactly matches any of the reference answers after normalization. The normalization process includes:
  • Lowercasing the text.
  • Removing punctuation and articles (a, an, the).
  • Replacing underscores with spaces.
  • Collapsing multiple spaces into a single space.
This ensures robustness to minor formatting or surface-level differences. A score of 1.0 is returned for an exact match; otherwise, 0.0.
C.2 F1 score calculation
The F1 score is computed at the token level based on normalized answers. For each reference answer, we calculate the precision and recall between the predicted and ground truth tokens, and derive the F1 score as their harmonic mean. The maximum F1 score across all ground truths is taken as the final score for a given prediction.
C.3 LLM-as-a-Judge evaluation
To evaluate semantic correctness that may not be captured by surface-level metrics, we adopt a GPT-based LLM-as-a-Judge (LJ) method. For each question, the model’s generated response and the reference answers are embedded into a prompt and passed to GPT-4o (version 2024-11-20). The model is instructed to answer "yes" if the response is semantically consistent with any of the correct answers (including paraphrases or synonyms), and "no" otherwise.
The prompt template is shown in Table 8:
Table 8: Prompt template used in LLM-as-a-Judge evaluation. The placeholders {correct_answer} and {model_response} are replaced at runtime.
*************Consider a knowledge Q&A RAG task to test the capability of a testing model, the correct answer list is:*************
{correct_answer}
*************Here is the model’s response:*************
{model_response}
*************Please check if the model’s answer is correct. As long as the model’s answer hits any item (or synonym) in the correct answer list, it can be considered correct. You only need to answer "yes" or "no".*************
When calling the GPT-4o API, we set the temperature to 0.0 to ensure deterministic judgment behavior.
Appendix D Details of Supervised Fine-Tuning
To highlight the benefits of reinforcement learning, we design two supervised fine-tuning (SFT) baselines:
SFT-direct.
This variant trains the model to generate the final answer directly. During training, we concatenate the input question with all retrieved paragraphs (in their original order) and instruct the model to output a short answer without any intermediate reasoning or evidence selection. This setup resembles conventional instruction-tuning on flat context.
SFT-reasoning.
This variant follows a two-stage pipeline to teach the model structured reasoning. Specifically, for each training instance, we collect:
  • Question $q$
  • Retrieved reference set $\mathcal{C}$
  • Ground-truth relevant reference IDs $\mathcal{I}$
  • Ground-truth final answer $\mathcal{O}$
These components are fed into the backbone model using a reasoning prompt shown in Table 9, prompting it to generate a coherent reasoning trace. The output reasoning trace is then integrated into the structured format described in Section 3.1:
$(q,\mathcal{C})\rightarrow(\mathcal{I},\mathcal{Z},\mathcal{O})$
This structured format is then used for supervised fine-tuning with full output supervision.
Table 9: Prompt format for collecting structured reasoning traces in SFT-reasoning. The model is instructed to derive the answer using only specified relevant references.
Generate a thinking process showing how to derive the given answer using ONLY the specified relevance IDs from provided references.
Question: {question}
References: {references}
Relevance IDs: {relevance_ids}
Answer: {answer}
Appendix E Details of Training Configuration
We train all models using the Open-R1 reinforcement learning framework with DeepSpeed ZeRO Stage 2 optimization and bfloat16 mixed-precision. The total batch size is 256. The learning rate is set to 3e-6. We generate 7 rollout samples per input for reward estimation. We set temperature = 0.9 during rollout, KL coefficient $\beta=0.04$, number of iterations per batch $\mu=1$, and clipping parameter $\epsilon=0.2$.
Appendix F KL Divergence Stabilization
To examine the effect of our stabilized KL divergence formulation (see Section 3.2), we compare its training dynamics with the standard KL estimator. As shown in Figure 4, the stabilized variant exhibits smooth and bounded updates throughout training, while the original estimator suffers from catastrophic spikes. This confirms that our modified KL term not only improves theoretical stability but also provides better empirical convergence in practice.
Figure 4: Training KL divergence curves for stabilized vs. unstabilized variants. ARENA’s stabilized KL (left) leads to smoother convergence, while the standard estimator (right) suffers from extreme spikes.
Appendix G Additional Experiemntal Results on Reward Variants
To further evaluate the impact of different accuracy reward designs during training, we conduct the same experiments on 2WikiMultiHopQA and MuSiQue. Results are shown in Appendix G and Appendix G. On 2WikiMultiHopQA, EM-based reward continues to yield the best performance across all three metrics, confirming its generalizability on structured reasoning tasks.
On MuSiQue, however, we observe that GPT- and F1-based judgment outperform EM-based reward. This may be attributed to the greater difficulty of the MuSiQue dataset—each question is paired with 20 retrieved paragraphs instead of 10—making exact matching harder and rendering fuzzy evaluation criteria more helpful for guiding learning. These findings suggest that while EM-based reward is generally more stable and interpretable, looser metrics might be preferable in high-noise or high-ambiguity retrieval settings.
Table 10: Reward method comparison on 2WikiMultiHopQA. EM-based reward remains the most effective.
Judge Method EM F1 LJ
GPT-4o 0.644 0.730 0.740
F1 0.646 0.736 0.732
EM (ours) 0.660 0.752 0.774
Table 11: Reward method comparison on MuSiQue. F1- and GPT-based rewards show advantage in this more challenging setting.
Judge Method EM F1 LJ
GPT-4o 0.430 0.533 0.516
F1 0.438 0.548 0.502
EM (ours) 0.400 0.520 0.508